# Personalized Attacks of Social Engineering in Multi-turn Conversations - LLM Agents for Simulation and Detection

### v1

@@ 5 Can LLM Simulations Proxy Realistic CSE Dynamics?
This section examines whether SE-VSim-generated SE conversations realistically capture real-world CSE attack dynamics. Instead of making definitive causal claims, we analyze observed attack success patterns as seen in Figure 3 and compare them with established psychological and cybersecurity research. By aligning our findings with prior studies on personality-based SE susceptibility, we assess whether SE-VSim can serve as reliable proxies for studying CSE threats.
Figure 3: Attack Success Distribution Over Personality Traits
@@ 5.1 Overall Correlation Between Personality Traits and Attack Outcomes
Conscientiousness correlates with increased vulnerability in compliance-driven interactions. Our analysis suggests that highly conscientious individuals are more susceptible to highly successful SE attacks when attackers simulate authority or professionalism. This finding is consistent with behavioral research on compliance Guadagno and Cialdini (2005); Schmeisser et al. (2021), which shows that conscientious individuals tend to adhere to perceived rules and norms, even when fabricated by an attacker. Real-world cybersecurity research further supports this pattern, indicating that individuals with high conscientiousness are more likely to follow directives without questioning their authenticity Rahman et al. (2022). Attackers exploiting corporate or bureaucratic environments frequently target conscientious individuals due to their strong adherence to procedural expectations.
Agreeable individuals are susceptible due to their trust-oriented nature. Figure 3 shows that high agreeableness is consistently linked to highly successful attacks. Agreeable individuals are often characterized by their desire to avoid conflict and maintain harmonious relationships Goldberg (2013); Cusack and Adedokun (2019), which can be exploited by attackers. In the context of SE attacks, these individuals may find it difficult to question or challenge requests, making them more likely to fall victim to tactics like phishing or pretexting, where attackers pose as trusted figures.
@@ 5.2 Fine-Grained Correlation by Attacker Role and Information Type
Professional contexts exacerbate vulnerability in highly conscientious individuals. In the Funding Agency scenario, individuals with high conscientiousness were significantly more vulnerable to highly successful SE attacks. This finding is grounded in the psychology of compliance Guadagno and Cialdini (2005); Schmeisser et al. (2021), particularly in formal or hierarchical environments, where individuals feel pressured to conform to perceived rules or expectations. In real-world SE attacks targeting corporate or financial environments, attackers often pose as authority figures (e.g., funding agencies, executives), knowing that highly conscientious individuals are more likely to comply with formal requests without questioning their authenticity.
Low agreeableness leads to successful trust-building by attackers in long-term engagements. In the same AF scenario, individuals with low agreeableness were less likely to disclose sensitive information immediately, yet they were often partially manipulated. This indicates that although these individuals resist initial engagement, attackers can still succeed in establishing trust over time. This is reflective of real-world attacks where attackers use prolonged approaches to gain trust.
@@ 5.3 Implications for Real-World SE Attack Prevention
Our analysis suggests that the simulated attack interactions capture well-documented behavioral vulnerabilities in SE. Particularly our findings demonstrate that personality traits, such as high conscientiousness and agreeableness, significantly influence the success of SE attacks. Studying how attackers exploit these traits offers a deeper understanding of victim behavior, revealing how psychological vulnerabilities are manipulated. This approach is essential for developing tailored SE defense strategies that can more effectively target individuals based on their specific traits.
6 Can We Defend against Realistic SE-attacks?
@@ 6.1.1 Baselines Detectors
We evaluate the performance of two categories of baseline detectors: (1) LLMs as Detectors, which include the Llama-3 8B model and the GPT-4 model in both zero-shot and few-shot settings, and (2) LLM-based Frameworks, which consist of the ConvoSentinel pipeline and our proposed approach. The LLM-based detectors are directly prompted to detect SE attacks in multi-turn conversations from our dataset, whereas the LLM-based frameworks employ additional reasoning mechanisms to enhance detection performance. For the few-shot baselines, each model is provided with one malicious and one benign example from a specific attack setting (AF, JO, or RE) to assess their ability to generalize across similar interactions.
6.1.2 Experiment Setting
In this evaluation, we split the malicious conversations into successful and partially successful attacks, using the successful label. Both the Llama-3 8B and GPT-4o-mini models are evaluated under zero-shot and few-shot configurations without any additional training or fine-tuning. The dataset split is described in Appendix E. The ConvoSentinel pipeline is run with its default configuration as described in the paper Ai et al. (2024), with the small adjustment of replacing the Llama 2 component of the pipeline with Llama 3. We also replaced the GPT-3.5 Turbo decision-making component of the ConvoSentinel pipeline with GPT-4o-mini and consequently saw marginal performance improvements. We employ standard metrics, F1-score and accuracy to measure the effectiveness of the models in identifying SE attempts, focusing on both fully and partially successful attacks.
@@ 6.1.3 Findings
As seen in Table 1 the performance of the evaluated detectors on our dataset remains suboptimal. Both the Llama-3 8B and GPT-4o-mini models, as well as the ConvoSentinel pipeline, exhibit low F1 and accuracy, particularly for multi-turn interactions lacking direct sensitive information exchange. When we restrict the analysis to fully successful SE attacks, where sensitive information is mostly disclosed, the models demonstrate improved performance. As shown in Figure 4, unsuccessful social engineering attempts are comparatively easier to detect, as the initiator often resorts to repetitive and overt requests for information, which clearly signal malicious intent. In contrast, partially successful scenarios present a greater challenge; these conversations often involve more subtle techniques, such as gradual trust-building and nuanced manipulation, rather than explicit information requests.
@@
Figure 4: Comparison of detection accuracy by success level of social engineering attempts, focusing on SE-OmniGuard and GPT-4o Few-shot detectors. Dashed lines represent overall accuracy for each detector
@@ 6.2.1 Framework Design
The control agent serves as the orchestrator of the detection process, conditioned by human expertise to assess the conversation based on factors such as sensitive information exchange, victim personality traits, and attacker strategies. The control agent delegates specific tasks to worker agents, each responsible for evaluating a particular aspect of the conversation. For instance, one worker agent focuses on the victim’s personality traits to detect if the attacker is exploiting any psychological vulnerabilities, while another worker agent analyzes the attack strategy to understand how the attacker manipulates the conversation. After each worker agent completes its task, the findings are reported to the control agent, which then synthesizes these insights to make a final decision on whether the conversation constitutes a malicious SE attempt. The prompts and implementation details of the SE-OmniGuard can be found in the appendix D. As a proof of concept, we assume that the framework has knowledge of the victim’s personality traits and operates with a predefined set of attack strategies. Improving the framework to infer personality traits dynamically and generalize to unseen attack strategies remains an area for future work.
@@ 6.2.3 Findings
The delegate-based detection framework significantly improves detection accuracy compared to existing LLM-based detectors as shown in Table 1. Notably, SE-OmniGuard performs well across all categories, demonstrating strong accuracy in detecting both unsuccessful and partially successful SE attacks as shown in Figure 4. The worker agents, by focusing on specific aspects of the conversation (such as personality traits or attack strategies), are more effective at detecting partially successful SE attacks, where the attacker builds trust without obtaining sensitive information immediately. The use of smaller LLMs as worker agents also results in significant cost savings, making the framework scalable for high-volume SE detection in multi-turn settings. Moreover, the control agent’s ability to synthesize the findings from the worker agents ensures that the final decision is well-informed and contextually grounded.
@@ 7 Conclusion
In this paper, we addressed the growing threat of multi-turn SE attacks facilitated by LLM agents. These attacks are more complex than single-instance interactions, and current detection methods often overlook partially successful SE attempts that involve trust-building without immediate sensitive information exchange. We introduced SE-VSim, an LLM-agentic framework designed to simulate SE attack mechanisms by generating realistic multi-turn conversations that account for victim personality traits and attack strategies. Based on insights from SE-VSim, we developed SE-OmniGuard, a proof-of-concept that uses a delegate design pattern, with a control agent and specialized worker agents analyzing specific conversation aspects, such as victim traits and attacker tactics. Our approach improves detection accuracy and cost-efficiency, making it scalable for real-world SE detection in high-volume conversational settings. The results demonstrate that SE-OmniGuard significantly outperforms existing detectors in identifying nuanced SE attacks, providing an effective and optimized direction for SE defense.
@@ 8.1 Constraints of the SE-VSim Simulation Framework
Additionally, the victim agents in SE-VSim are modeled with varying psychological profiles based on the Big Five personality traits. While the psychology community generally agrees on the usefulness of these traits for predicting deception, there is no assurance that the profiles generated comprehensively cover all relevant personality types. Finally, SE-VSim is LLM-generated and thus prone to hallucination and sycophancy, introducing potential misrepresentation of real-world CSE attacks. Despite these limitations, SE-VSim is the first dataset of its kind and could be expanded to include additional attack situations and personality types in the future.
@@ 10 Acknowledgments
This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA). The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
Appendix A SE-VSim Implementation Details
A.1 SE-VSim Parameters
@@ A.2 Prompting
The Table 5 summarizes the values used for the attacker’s placeholders in different attack scenarios.
@@
A.3 Personality Trait Descriptions
@@
Table 6: Personality Traits and Personas
Appendix B Annotation Guidelines
@@ B.1 Human Annotation Guidelines
Annotation Process: Each annotator will be provided with a unique spreadsheet for recording their annotations. An Annotation Spreadsheet is provided as an example. Using this template, make a working copy with the annotator name added to the title.
@@ B.2 LLM-Judge Annotations
The prompt used for the LLM-judge is detailed in Table 7. To enhance the clarity and effectiveness of the LLM-judge’s decisions, the prompt includes set of examples extracted from frequently asked questions (FAQs) and clarifications exchanged between the authors and human annotators during the annotation process. This additional context was designed to provide the LLM-judge with a more comprehensive understanding of the annotation guidelines and improve the quality of its annotations.
@@
Appendix C Attack Strategy Annotation Details
@@
C.1 Attack Strategies and Sub-Categories
@@
  • Source Credibility and Authority: People tend to comply with requests from perceived authority figures. Symbols of authority like uniforms or badges can increase compliance.
@@
C.2 LLM-Based Annotation Methodology
Due to the complexity and multilabel nature of this task, manual annotation by human annotators would have been prohibitively costly. Instead, we employed an LLM-judge GPT-4o-mini (gpt-4o-mini-2024-07-18) to annotate the conversations. The LLM performed the following steps:
@@ C.2 LLM-Based Annotation Methodology
C.3 Attack Strategy Analysis
Our analysis of attack strategies reveals that attacker agent tailor its strategies based on the victim’s personality traits and the attack context. For instance, agreeable and conscientious individuals are particularly vulnerable to tactics that rely on trust and perceived authority, while extraverts and less conscientious individuals are more likely to engage with strategies that emphasize urgency and social cues. The analysis of these specific attack scenarios highlights the importance of contextual factors in SE attacks. The success of an attack is not only determined by the victim’s personality traits but also by the nature of the interaction and the perceived authority or formality of the context.
@@ C.3 Attack Strategy Analysis
C.4 Attack Strategy Analysis
@@
Persuasion and social proof strategies are most effective against highly agreeable and conscientious individuals By analyzing the attack strategies used by the attacker agent, we observe that persuasion and social proof tactics are particularly effective against individuals with high agreeableness and conscientiousness. These personality types are more likely to comply with social cues or requests from authority figures, making them prime targets for attacks that rely on establishing trust and legitimacy.
@@
Appendix D SE-OmniGuard Implementation Details
This section provides the prompts used for the worker agents and the control agent in the SE-OmniGuard framework, along with the generation parameters used for the LLM-based analysis. Prompt used for control agent is noted in Table 10 and worker agent prompts can be found in Table 9.
@@ D.1 Generation Parameters
Appendix E Data Construction for Experiments
E.1 Data Split

### v2

@@ 5 Can LLM Simulations Proxy Realistic CSE Dynamics?
This section examines whether SE-VSim-generated SE conversations capture real-world CSE attack dynamics. Instead of making definitive causal claims, we analyze observed attack success patterns as seen in Figure 2(b) and compare them with established psychological and cybersecurity research. By aligning our findings with prior studies on personality-based SE susceptibility, we assess whether SE-VSim can serve as a reliable proxy for studying CSE threats.
@@ 5.1 Overall Correlation Between Personality Traits and Attack Outcomes
Conscientiousness correlates with increased vulnerability in compliance-driven interactions. Our analysis suggests that highly conscientious individuals are more susceptible to highly successful SE attacks when attackers simulate authority or professionalism. This finding is consistent with behavioral research on compliance Guadagno & Cialdini (2005); Schmeisser et al. (2021), which shows that conscientious individuals tend to adhere to perceived rules and norms, even when fabricated by an attacker. Real-world cybersecurity research further supports this pattern, indicating that individuals with high conscientiousness are more likely to follow directives without questioning their authenticity Rahman et al. (2022). Attackers exploiting corporate or bureaucratic environments frequently target conscientious individuals because they adhere strongly to procedural expectations.
Agreeable individuals are susceptible due to their trust-oriented nature. Figure 2(b) shows that high agreeableness is consistently linked to highly successful attacks. Agreeable individuals are often characterized by their desire to avoid conflict and maintain harmonious relationships Goldberg (2013); Cusack & Adedokun (2019), which attackers can exploit. In the context of SE attacks, these individuals may find it difficult to question or challenge requests, making them more likely to fall victim to tactics like phishing or pretexting, where attackers pose as trusted figures.
@@ 5.2 Fine-Grained Correlation by Attacker Role and Information Type
Professional contexts exacerbate vulnerability in highly conscientious individuals. In the Funding Agency scenario, individuals with high conscientiousness were significantly more vulnerable to highly successful SE attacks. This finding is grounded in the psychology of compliance Guadagno & Cialdini (2005); Schmeisser et al. (2021), particularly in formal or hierarchical environments, where individuals feel pressured to conform to perceived rules or expectations. In real-world SE attacks targeting corporate or financial environments, attackers often pose as authority figures (e.g., funding agencies, executives), knowing that highly conscientious individuals are more likely to comply with formal requests without questioning their authenticity.
Low agreeableness leads to successful trust-building by attackers in long-term engagements. In the same AF scenario, individuals with low agreeableness were less likely to disclose sensitive information immediately, yet they were often partially manipulated. This indicates that although these individuals resist initial engagement, attackers can still succeed in establishing trust over time. It is also reflective of real-world attacks where attackers use prolonged approaches to gain trust.
@@ 5.3 Implications for Real-World SE Attack Prevention
Our analysis suggests that the simulated attack interactions capture well-documented behavioral vulnerabilities in SE. Notably, our findings demonstrate that personality traits, such as high conscientiousness and agreeableness, significantly influence the success of SE attacks. Studying how attackers exploit these traits offers a deeper understanding of victim behavior, revealing how psychological vulnerabilities are manipulated. This approach is essential for developing tailored SE defense strategies that can more effectively target individuals based on their specific traits.
6 Can We Defend Against Realistic SE Attacks?
@@ 6.1.1 Baselines Detectors
We evaluate the performance of two categories of baseline detectors: (1) LLM-based Detectors, which include the Llama-3 8B model and the GPT-4 model in both zero-shot and few-shot settings, and (2) LLM-based Frameworks, which consist of the ConvoSentinel pipeline and our proposed approach. The LLM-based detectors are directly prompted to detect SE attacks in multi-turn conversations from our dataset, whereas the LLM-based frameworks employ additional reasoning mechanisms to enhance detection performance. For the few-shot baselines, each model is provided with one malicious and one benign example from a specific attack setting (AF, JO, or RE) to assess their ability to generalize across similar interactions.
6.1.2 Experiment Settings
In this evaluation, we split the malicious conversations into successful and partially successful attacks, using the successful label. Both the Llama-3 8B and GPT-4o-mini models are evaluated under zero-shot and few-shot configurations without any additional training or fine-tuning. The dataset split is described in Appendix F. The ConvoSentinel pipeline is run with its default configuration as described in the paper Ai et al. (2024), with the small adjustment of replacing the Llama 2 component of the pipeline with Llama 3. We also replaced the GPT-3.5 Turbo decision-making component of the ConvoSentinel pipeline with GPT-4o-mini and consequently saw marginal performance improvements. We employ standard metrics, F1-score, and accuracy to measure the effectiveness of the models in identifying SE attempts, focusing on both fully and partially successful attacks.
@@ 6.1.3 Findings
As seen in Table 1, the performance of the evaluated detectors on our dataset remains suboptimal. Both the Llama-3 8B and GPT-4o-mini models, as well as the ConvoSentinel pipeline, exhibit low F1 and accuracy, particularly for multi-turn interactions lacking direct sensitive information exchange. When we restrict the analysis to fully successful SE attacks, where sensitive information is mostly disclosed, the models demonstrate improved performance. As shown in Figure 3, unsuccessful social engineering attempts are comparatively easier to detect, as the initiator often resorts to repetitive and overt requests for information, which clearly signal malicious intent. In contrast, partially successful scenarios present a greater challenge; these conversations often involve more subtle techniques, such as gradual trust-building and nuanced manipulation, rather than explicit information requests.
@@
Figure 3: Comparison of detection accuracy by the success level of social engineering attempts, focusing on SE-OmniGuard and GPT-4o few-shot detectors. Dashed lines represent the overall accuracy for each detector.
@@ 6.2.1 Framework Design
The control agent serves as the orchestrator of the detection process, conditioned by human expertise to assess the conversation based on factors such as sensitive information exchange, victim personality traits, and attacker strategies. The control agent delegates specific tasks to worker agents, each responsible for evaluating a particular aspect of the conversation. For instance, one worker agent focuses on the victim’s personality traits to detect if the attacker is exploiting any psychological vulnerabilities, while another worker agent analyzes the attack strategy to understand how the attacker manipulates the conversation. After each worker agent completes its task, the findings are reported to the control agent, which then synthesizes these insights to make a final decision on whether the conversation constitutes a malicious SE attempt. The prompts and implementation details of the SE-OmniGuard are included in Appendix E. As a proof of concept, we assume that the framework has knowledge of the victim’s personality traits and operates with a predefined set of attack strategies. Improving the framework to infer personality traits dynamically and generalize to unseen attack strategies remains an area for future work.
@@ 6.2.3 Findings
The delegate-based detection framework significantly improves detection accuracy compared to existing LLM-based detectors, as shown in Table 1. Notably, SE-OmniGuard performs well across all categories, demonstrating strong accuracy in detecting both unsuccessful and partially successful SE attacks as shown in Figure 3. The worker agents, by focusing on specific aspects of the conversation (such as personality traits or attack strategies), are more effective at detecting partially successful SE attacks, where the attacker builds trust without obtaining sensitive information immediately. The use of smaller LLMs as worker agents also results in significant cost savings, making the framework scalable for high-volume SE detection in multi-turn settings. Moreover, the control agent’s ability to synthesize the findings from the worker agents ensures that the final decision is well-informed and contextually grounded.
@@ 7 Conclusion
In this paper, we addressed the growing threat of multi-turn SE attacks facilitated by LLM agents. These attacks are more complex than single-instance interactions, and current detection methods often overlook partially successful SE attempts that involve trust-building without immediate sensitive information exchange. We introduced SE-VSim, an LLM-agentic framework designed to simulate SE attack mechanisms by generating realistic multi-turn conversations that account for victim personality traits and attack strategies. Based on insights from SE-VSim, we developed SE-OmniGuard, a proof of concept that uses a delegate design pattern, with a control agent and specialized worker agents analyzing specific conversation aspects, such as victim traits and attacker tactics. Our approach improves detection accuracy and cost-efficiency, making it scalable for real-world SE detection in high-volume conversational settings. The results demonstrate that SE-OmniGuard significantly outperforms existing detectors in identifying nuanced SE attacks, providing an effective and optimized direction for SE defense.
@@ 8.1 Constraints of the SE-VSim Simulation Framework
Additionally, the victim agents in SE-VSim are modeled with varying psychological profiles based on the Big Five personality traits. While the psychology community generally agrees on the usefulness of these traits for predicting deception, there is no assurance that the profiles generated comprehensively cover all relevant personality types. Finally, SE-VSim is LLM-generated and thus prone to hallucination and sycophancy, which may lead to the misrepresentation of real-world CSE attacks. Despite these limitations, SE-VSim is the first dataset of its kind and could be expanded to include additional attack situations and personality types in the future.
@@ 10 Acknowledgments
This material is based upon work supported by the Defense Advanced Research Projects Agency under Contract No. HR001120C0123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA).
Appendix A Appendix
Category Method Accuracy F1
AF JO RE Overall AF JO RE Overall
Llama 3AF Few-Shot 0.615 0.580 0.585 0.593 0.709 0.693 0.691 0.698
Llama 3JO Few-Shot 0.590 0.590 0.515 0.565 0.667 0.682 0.625 0.658
Llama 3RE Few-Shot 0.630 0.575 0.535 0.580 0.713 0.679 0.646 0.679
GPT-4oZero-Shot 0.615 0.520 0.600 0.578 0.374 0.077 0.333 0.271
GPT-4oAF Few-Shot 0.680 0.600 0.745 0.675 0.543 0.333 0.657 0.523
GPT-4oJO Few-Shot 0.715 0.605 0.720 0.680 0.612 0.347 0.616 0.536
GPT-4oRE Few-Shot 0.670 0.580 0.745 0.665 0.522 0.276 0.662 0.504
ConvoSentinelGPT-4o 0.710 0.610 0.690 0.680 0.710 0.540 0.660 0.640
SE-OmniGuard (Ours) 0.740 0.835 0.865 0.813 0.775 0.814 0.862 0.815
Table 2: Baseline ablation and generalization: few-shot baselines were trained using only one data class (AF, JO, or RE) and evaluated on both seen and unseen data classes.
Appendix B SE-VSim Implementation Details
B.1 SE-VSim Parameters
@@ B.2 Prompting
Table 5 summarizes the values used for the attacker’s placeholders in different attack scenarios.
@@
B.3 Personality Trait Descriptions
@@
Table 6: Personality Traits and Personas.
Appendix C Annotation Guidelines
@@ C.1 Human Annotations
Annotation Process: Each annotator will be provided with a unique spreadsheet for recording their annotations. An annotation Spreadsheet is provided as an example. Using this template, make a working copy with the annotator’s name added to the title.
@@ C.2 LLM-Judge Annotations
The prompt used for the LLM-judge is detailed in Table 7. To enhance the clarity and effectiveness of the LLM-judge’s decisions, the prompt includes a set of examples extracted from frequently asked questions (FAQs) and clarifications exchanged between the authors and human annotators during the annotation process. This additional context was designed to provide the LLM-judge with a more comprehensive understanding of the annotation guidelines and improve the quality of its annotations.
@@
Appendix D Attack Strategy Annotation Details
@@
D.1 Attack Strategies and Sub-Categories
@@
  • Source Credibility and Authority: People tend to comply with requests from perceived authority figures. Symbols of authority, like uniforms or badges, can increase compliance.
@@
D.2 LLM-Based Annotation Methodology
Due to its complexity and multilabel nature, manual annotation by human annotators would have been prohibitively costly. Instead, we employed an LLM-judge GPT-4o-mini (gpt-4o-mini-2024-07-18) to annotate the conversations. The LLM performed the following steps:
@@ D.2 LLM-Based Annotation Methodology
D.3 Attack Strategy Analysis
Our analysis of attack strategies reveals that attacker agents tailor their strategies based on the victim’s personality traits and the attack context. For instance, agreeable and conscientious individuals are particularly vulnerable to tactics that rely on trust and perceived authority, while extraverts and less conscientious individuals are more likely to engage with strategies that emphasize urgency and social cues. The analysis of these specific attack scenarios highlights the importance of contextual factors in SE attacks. The success of an attack is not only determined by the victim’s personality traits but also by the nature of the interaction and the perceived authority or formality of the context.
@@ D.3 Attack Strategy Analysis
D.4 Attack Strategy Analysis
@@
Persuasion and social proof strategies are most effective against highly agreeable and conscientious individuals. By analyzing the attack strategies used by the attacker agent, we observe that persuasion and social proof tactics are particularly effective against individuals with high agreeableness and conscientiousness. These personality types are more likely to comply with social cues or requests from authority figures, making them prime targets for attacks that rely on establishing trust and legitimacy.
@@
Appendix E SE-OmniGuard Implementation Details
This section provides the prompts used for the worker agents and the control agent in the SE-OmniGuard framework, along with the generation parameters used for the LLM-based analysis. The prompt used for the control agent is noted in Table 10 , and worker agent prompts are in Table 9.
@@ E.1 Generation Parameters
Appendix F Data Construction for Experiments
F.1 Data Split